From GlueXWiki
Jump to: navigation, search

May 16, 2013 FDC meeting


  1. Production Construction Tracking
    • Spare package status
  2. Installation (Lubomir, Bill)
    • Update on the mock-up
  3. Engineering (Bill)
    • Cooling system status
  4. Electronics (Chris, Nick)
  5. Full electronics tests FDC E-log (Beni)
  6. Other


Participants: Casey, Bill, Dave, Chris, Simon, Beni, and Lubomir.


- Spare package: interestingly the second cell with wire frame #6 showed the same problems as with wire frame #3 before: no signals on chan.1, point source of noise in one of the middle wires and glow discharge (sudden increase of the current on both sense and field wires). The difference is that the source of the noise was at the end of the wire (as picked up by the strips), and also that the glow discharge disappears when removing some of the cathode cards. All this was pointing that the problems are caused by the tip of the wire sticking out of the solder. As for chan.1 we looked at the previous wire frame #3 and found the HV resistor for wire #1 had a crack. On Tuesday we opened the top cell (with the help of Mark Bear) and found exactly what we expected. The cracked resistor was replaced, the solder pad re-flowed to cover the wire, and then we closed the package. Will start testing it on Friday afternoon; Dave will bring the NIM power supply that we share with the magnet people. The two wire frames #3 and #6 were done at the same time by the same people using same tools, obviously this explains the similarities in their problems.

- At the beginning of next week (most likely Tuesday) we will install two new cells (wire frames #27 and #28). The last two cells will have thicker spacers. We discussed what kind of holes we want to drill on these spacer rings to have the same pressure in the cells. The best is to have separate flow controller for the last two cells and to regulate the flow so that there's no big pressure difference between the cells. Bill propose to use a simple plastic part used in fans to spread the air. It will be mounted inside the spacer rings.


- Mock-up: Casey adjusted and connected all the eight rails of the two carts at Blue Crab, now we can continue with the test putting the packages on the cart. Before that (next Wednesday) Casey with the help of Bill will make one of the inter-package spacers to install it between two of the packages.


- Bill: there are problems with the water chiller; Mike Sprouse suggested several things: to measure the temperature at the input and output of the chiller and estimate the heat produced by the chiller itself. He suggested also to clean the heat exchanger. Another option is to find another water chiller. After the meeting Bill with Casey found that the water chiller itself is producing a lot of heat, therefore Bill borrowed another chiller from the accelerator people. So, the plan now is to replace the water chiller which will happen at the beginning of next week.

Full electronics tests

- Beni: no news for the firmware of fADC125. The plan is to move the third package on Monday. If the firmware/software is not ready by the end of next week we will get the modules from Cody as they are and read out only the old modules.